OpenStudy (anonymous):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 18 Overlapping lines Parallel lines Intersecting lines

OpenStudy (anonymous):

Think this is same line

OpenStudy (anonymous):

or overlapping lines if you multiply the first equation by three

OpenStudy (anonymous):

overlapping :D

OpenStudy (anonymous):

correction parallel lines

OpenStudy (anonymous):

You think it's parallel lines @JAMG

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x + y = 6 y = -2x + 6 6x + 3y = 18 3y = -6x + 18 y = -6/3 + 6 y = -2x + 6 same slope...parallel lines

OpenStudy (anonymous):

oh sorry i thought it was parallel has no solutions :( oops

OpenStudy (anonymous):

Johnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school? 22 42 51 60

OpenStudy (anonymous):

can you clarify this one kellie i think its 22 boys

OpenStudy (anonymous):

22*2 +16 + 22 = 82 44+16+22=82 60+22=82

OpenStudy (anonymous):

its 60 then right @JAMG

OpenStudy (anonymous):

22 is correct I just showed you that it makes sense from the problem

OpenStudy (anonymous):

and yes there are 60 boys

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

b + g = 82 b = 2g + 16 2g + 16 + g = 82 3g = 82 - 16 3g = 66 g = 22 b + g = 82 b + 22 = 82 b = 82 - 22 b = 60 check.. b = 2g + 16 60 = 2(22) + 16 60 = 44 + 16 60 = 60 (correct) there are 60 boys and 22 girls

OpenStudy (anonymous):

sorry, im not at my best right now :(

OpenStudy (anonymous):

don't worry about it...if I had $1 for every mistake I made, I would be rich

OpenStudy (anonymous):

loll, thanks @JAMG and @kelliegirl33 also no worries jamg :)

OpenStudy (anonymous):

glad to help :)